OpenStudy (anonymous):

The ratio of trumpet players to trombone players in the band is 5:3. What does this mean? For every 5 trumpet players, there are 3 trombone players. There are 2 more trumpet players than trombone players. There are 5 more trumpet players than trombone players. For every 5 trombone players, there are 3 trumpet players.

OpenStudy (anonymous):

hi you got here fast

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

or d

OpenStudy (mathstudent55):

Without knowing the number of players in the band, we know the ratio of trumpet players to trombone players is 5 to 3. For example, there could be 5 trumpet players and 3 trombone players, or 10 trumpet players and 6 trombone players, or 15 trumpet players and 9 trumpet players, etc.

OpenStudy (mathstudent55):

Look closely at d. It shows more trombone players than trumpet players. That is not what the ratio tells us.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

d is out.

OpenStudy (mathstudent55):

Now look at a. According to the ratio, for every 5 trumpet players there are 3 trombone players. That is exactly what choice a states.

OpenStudy (mathstudent55):

Choice b is true only if there are exactly 5 trumpet players and 3 trombone players, but we are not told the exact number of players. We only know the ratio.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

and c are left

OpenStudy (mathstudent55):

Now look at choice c. Choice c states there are 5 more trombone players than trumpet players. Is this possible? If there are 5 trombone players and 3 trumpet players, there are 2 more trombone players that trumpet players. If there are 10 trombone players and 6 trumpet players, there are 4 more trombone players that trumpet players. If there are 15 trombone players and 9 trumpet players, there are 6 more trombone players that trumpet players. The number of trombone players can be 2, 4, 6, 8, 10, etc. more than trombone players, but it can never be 5 more.

OpenStudy (mathstudent55):

a is the only choice that is true for any number of players, so a is the answer.

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

Betty and Adam are playing cards. Betty has won 5 hands and Adam has won 20 hands. What is the ratio of the number of hands Betty has won to the number of hands Adam has won? 1:4 4:1 5:1 1:5

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Just using the actual number of hands each one won, what is the ratio?

OpenStudy (anonymous):

um wait let me see

OpenStudy (anonymous):

is the answer a

OpenStudy (anonymous):

1;4

OpenStudy (mathstudent55):

Yes. Betty won 5. Adam won 20 The ratio of wins of Betty to Adam is 5:20 5:20 can also be written as a fraction, 5/20 5/20 = 1/4 = 1:4

OpenStudy (anonymous):

ok